40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1439     Entry time: Sun Mar 29 13:44:27 2009
Author: steve 
Type: Update 
Category: SUS 
Subject: ETMY sus damping restored 

ETMY sus damping was found to be tripped.

It was retored.

All fluorecent light were turned off. Please try to conserve some energy.

ELOG V3.1.3-